# LogPrism: Unifying Structure and Variable Encoding for Effective Log Compression

### v1

@@ IV-E RQ4: Robustness Analysis: The Impact of Data Granularity
Compression Ratio Analysis: In single-archive mode, LogPrism achieves a 278.40% increase in average CR over its own chunked performance, with the gain on the Windows dataset reaching a massive 755.33%. When compared to Denum operating in the same global configuration, LogPrism outperforms it on 13 of the 16 datasets. The improvements are particularly significant on complex logs such as Windows (+22.49%) and Thunderbird (+21.29%). On two of the remaining three datasets (Hadoop, HDFS), LogPrism remains highly competitive, following Denum with a small difference of less than 2.50%. The only notable exception is the Spark dataset, where Denum’s specialized handling of specific numeric sequences proves more efficient. Overall, these results validate that LogPrism scales effectively to leverage global context to achieve superior compression.
@@ V Related Work
V-A General-purpose Compression Algorithms
General-purpose compressors provide a baseline for evaluating specialized methods. These algorithms can be categorized into three primary families. Dictionary-based approaches, exemplified by LZMA, achieve compression by replacing repeated byte sequences with references to a dictionary. Prediction-based methods, such as PPMd, leverage statistical models to encode characters based on their preceding context. Block-sorting algorithms, notably bzip2, utilize the Burrows-Wheeler Transform to cluster similar characters, thereby enhancing subsequent encoding efficiency.
@@ V-A General-purpose Compression Algorithms
While these methods exhibit strong performance on generic text data, they fundamentally operate on undifferentiated byte streams without awareness of log structure. Consequently, their effectiveness diminishes significantly when confronted with logs containing high-entropy variables such as unique request IDs and dynamic numerical values, as they cannot exploit the inherent semi-structured redundancy that spans log entries.
V-B Log-Specific Compression Methods
V-B1 Parser-Based Approaches
The parser-based paradigm, which represents the dominant methodology in contemporary log compression research, follows a two-stage workflow. In the first stage, a log parser analyzes the input to extract fixed templates and separate variable parameters. Templates are typically inferred from statistical analysis of a representative log sample. In the second stage, these components are compressed independently, with template identifiers being dictionarized and parameter sequences processed separately.
This paradigm has undergone continuous refinement. LogZip [23] established the foundational framework, utilizing the Drain parser combined with an iterative clustering algorithm for pattern mining. LogReducer [33] advanced the field by introducing specialized techniques for numerical parameters, including delta encoding for timestamps, correlation identification, and elastic encoding schemes. LogShrink [21] extended this analysis to encompass all parameter types, employing longest common subsequence techniques and entropy-based analysis to identify commonality and variability patterns. Beyond compression effectiveness, several systems have integrated query capabilities into the compressed representation. CLP [29] and LogGrep [32] enable efficient search operations directly on compressed log data, demonstrating the potential for compression methods to support downstream analytical tasks.
However, the parser-based paradigm carries an inherent limitation: the rigid separation of templates and parameters occurs early in the processing pipeline. This irreversible categorization can destroy contextual correlations that span structural and variable tokens, potentially limiting compression effectiveness when these cross-boundary patterns dominate the redundancy structure.
V-B2 Non-Parser-Based Approaches
Non-parser-based methods eschew explicit structural extraction, instead applying unified compression strategies to entire log lines. Early work in this category includes LogArchive [7], which groups similar logs using similarity functions, and MLC [12], which identifies redundancy through block-level deduplication. The most recent advancement in this category is the number-centric paradigm, exemplified by Denum [38]. Rather than parsing logs into templates and parameters, Denum performs a global binary classification of all tokens into numeric and non-numeric categories. It then applies specialized compression techniques optimized for each category, leveraging the arithmetic properties of numeric sequences while handling non-numeric strings as independent streams.
While this approach avoids the template inference overhead of parser-based methods, it still performs an early, global token categorization. This design choice can lead to the loss of correlation information that crosses the numeric/non-numeric boundary, particularly when the redundancy structure involves co-occurring patterns of both types.
Our work introduces a fundamentally different paradigm that unifies structural extraction and pattern encoding through hierarchical redundancy mining. Rather than pre-committing to fixed categories, our approach progressively distills log data through multiple stages, prioritizing the encoding of high-frequency aggregate patterns that preserve full contextual information. This allows the compression of entire “structure+variable” sequences with single identifiers, exploiting deep redundancies in log data.

### v2

@@ IV-E RQ4: Robustness Analysis: The Impact of Data Granularity
Compression Ratio Analysis: In single-archive mode, LogPrism achieves a 273.27% increase in average CR over its own chunked performance, with the gain on the Windows dataset reaching a massive 739.37%. When compared to Denum operating in the same global configuration, LogPrism outperforms it on 13 of the 16 datasets. The improvements are particularly significant on complex logs such as Windows (+22.49%) and Thunderbird (+21.29%). On two of the remaining three datasets (Hadoop, HDFS), LogPrism remains highly competitive, following Denum with a small difference of less than 2.50%. The only notable exception is the Spark dataset, where Denum’s specialized handling of specific numeric sequences proves more efficient. Overall, these results validate that LogPrism scales effectively to leverage global context to achieve superior compression.
@@ V Related Work
This section discusses the landscape of compression techniques relevant to log data, organized by their foundational design philosophies.
V-A General-purpose Compression Approaches
@@
These algorithms eliminate statistical redundancy in data for compression, which can be categorized into three primary families. Dictionary-based approaches, exemplified by LZMA, achieve compression by replacing repeated byte sequences with references to a dictionary. Prediction-based methods, such as PPMd, leverage statistical models to encode characters based on their preceding context. Block-sorting algorithms, notably bzip2, utilize the Burrows-Wheeler Transform to cluster similar characters, thereby enhancing subsequent encoding efficiency. While these methods exhibit strong performance on generic text data, they fundamentally operate on undifferentiated byte streams without awareness of log structure. Consequently, their effectiveness diminishes significantly when confronted with logs containing high-entropy variables such as unique request IDs and dynamic numerical values, as they cannot exploit the inherent semi-structured redundancy that spans log entries.
V-B Log-specific Compression Approaches
The dominant methodology in log compression is the parser-based paradigm [23, 33, 21, 29, 32], which decouples structure extraction from variable encoding. LogZip established this framework by combining the Drain parser with iterative clustering. Subsequent works focused on optimizing the separated components: LogReducer introduced delta encoding for timestamps, while LogShrink utilized entropy-based analysis to identify variable patterns. Additionally, systems like CLP and LogGrep have extended this paradigm to enable direct search on compressed data. To avoid the overhead of explicit template extraction, alternative approaches have emerged. Early works, such as LogArchive [7] and MLC [12], applied similarity grouping or block-level deduplication to compress logs. More recently, the number-centric paradigm, exemplified by Denum [38], performs a global binary classification to split tokens into numeric and non-numeric streams.
However, state-of-the-art methods share a fundamental limitation: the reliance on early, rigid token categorization. Parser-based methods separate templates from variables, while number-centric methods separate numerics from strings. This irreversible decoupling destroys contextual correlations that span these boundaries. Instead, LogPrism unifies structural extraction and pattern encoding to mine deep “structure+variable” redundancies that existing methods ignore.